### v1

# Range Image-Based Implicit Neural Compression for LiDAR Point Clouds

Abstract
This paper presents a novel scheme to efficiently compress Light Detection and Ranging (LiDAR) point clouds, enabling high-precision 3D scene archives, and such archives pave the way for a detailed understanding of the corresponding 3D scenes. We focus on 2D range images (RIs) as a lightweight format for representing 3D LiDAR observations. Although conventional image compression techniques can be adapted to improve compression efficiency for RIs, their practical performance is expected to be limited due to differences in bit precision and the distinct pixel value distribution characteristics between natural images and RIs. We propose a novel implicit neural representation (INR)–based RI compression method that effectively handles floating-point valued pixels. The proposed method divides RIs into depth and mask images and compresses them using patch-wise and pixel-wise INR architectures with model pruning and quantization, respectively. Experiments on the KITTI dataset show that the proposed method outperforms existing image, point cloud, RI, and INR-based compression methods in terms of 3D reconstruction and detection quality at low bitrates and decoding latency.
Index Terms:
LiDAR, Point Clouds, Range Image INR.
I Introduction
Light Detection and Ranging (LiDAR) sensors have gained significant attention not only in online applications but also in offline applications. In such offline applications, memory-efficient and precise three-dimensional (3D) scenes should be stored in advance and the 3D scenes should be smoothly retrieved from the storage based on the user demand for applications of 3D scene understanding such as digital archiving, environmental monitoring, navigation, and geological surveying  [1, 2, 3]. LiDAR sensors scan the physical space with the ego-centric coordinate and measure the distance to the closest point on surrounding objects for each angle, allowing the creation of a point cloud with 3D points corresponding to the intersection of laser beams with objects ahead. As the resolution of LiDAR sensors increases, effectively storing and transmitting LiDAR scans becomes a significant challenge, primarily due to the substantial volume of each LiDAR sequence.
Although LiDAR scans are typically represented as 3D point clouds, they can also be expressed as a single-channel image, referred to as a two-dimensional (2D) range image (RI) [4], where point clouds captured in an ego-centric spherical coordinate system are projected onto a panoramic image. The x-y axes of the RI image correspond to the azimuth and elevation angles in the 3D spherical coordinate system, while each pixel value represents the distance to the corresponding point in that direction. Whereas 3D point clouds require $3N$ values to represent the locations of $N$ point measurements, RI s require only $N$ pixels at most, thus demonstrating their compactness.
We can pursue methods to further compress RI s. One potential solution is to adapt conventional lossy image compression techniques, such as Joint Photographic Experts Group (JPEG) [5] and Joint Photographic Experts Group 2000 (JPEG2000) [6]. However, these methods are based on integer precision for pixel values, which is incompatible for RI s whose pixels are represented by single or double precision floating-point numbers to precisely express the distance to the corresponding point measurements. We can still utilize these compression methods by adjusting the bit precision of RI s to align with them, but this naturally leads to degradation in 3D point cloud reconstruction performance due to inadequate distance resolution. Another drawback of conventional compression methods lies in their strategy: they use block-based discrete cosine transform (DCT) and apply coarse quantization to high-frequency components based on human visual perception. This approach leads to significant degradation in the decoding performance of RI s, as RI s are characterized by large and sudden changes in the pixel value between foreground objects and distant backgrounds or pixels without any assigned point measurement.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
To effectively compress RI s while preserving their fine precision and high-frequency changes in pixel value, this paper presents a novel RI compression method inspired by implicit neural representation (INR)-based image compression technique [7]. INR [8, 9] is a lightweight representation of multidimensional signals by compressing them into shallow neural networks (NNs). Specifically, INR overfits NNs with a limited number of parameters to the signals of interest through supervised learning, and the trained parameters become the compressed signal representation by providing the mapping function from signal indices, for example, coordinates on the image plane, to the corresponding signals. A primary challenge for INR-based signal compression is to ensure the precision of high-frequency details while simultaneously managing the constraints imposed by the limited model size. To address this challenge, we propose an extended INR training approach that incorporates both a mask image and an RI for point depth information. The mask image is a binary map that indicates whether each pixel corresponds to a projected 3D point (zero) or not (one). We begin by generating a mask image from the RI, followed by learning two separate coordinate-to-value mappings using distinct INR architectures: one for depth INR and one for the mask INR. During decoding, these trained INR s reconstruct both the depth and mask images, and the final reconstructed RI is obtained by applying the mask to the depth image. Although the reconstructed depth image may contain values for pixels that do not correspond to any 3D points, applying the mask enforces hard thresholding, effectively removing these artifacts. This process ensures a high-quality reconstructed RI, with sharp edges accurately preserved.
The contributions of our study are three-fold:
  • To the best of our knowledge, this is the first paper to propose an INR-based intra RI compression method specifically designed for LiDAR measurements projected onto high-precision floating-point images.
  • We extend the INR compression approach to incorporate the mask INR that explicitly represents whether any point measurements are assigned to each pixel on the RI or not, allowing efficient elimination of false point estimation on reconstructed depth images in the decoding process.
  • We evaluate our proposed method using KITTI dataset [10] and compare its performance with existing baselines including conventional image compression, point cloud compression (PCC) [11], and RI-based and INR-based image compression, and show the better rate-distortion (R-D) performance and downstream task quality of our method than the baselines.
II Related Work
II-A Point Cloud Compression
The measured distance from LiDAR sensors is usually represented as a 3D point cloud. Each point cloud consists of a set of 3D points, and each point is defined by 3D coordinates, i.e., (X, Y, Z). The graph-based and tree-based compression methods have been proposed to compress coordinate information, that is, geometry information. The graph-based methods regard the 3D points as graph signals and define graph Fourier transform (GFT) for frequency conversion in the graph domain. [12, 13, 14] utilized GFT for the geometry compression. Other studies [15, 16] reduce the storage and transmission costs for graph signal reconstruction. The tree-based compression method is another popular strategy for compressing geometry information. The typical way is octree-based representation, such as point cloud library (PCL) and geometry-based point cloud compression (G-PCC) [17, 11]. Some recent studies have been proposed to improve the efficiency of geometry compression using traditional signal processing [18] and deep neural network (DNN) [19, 20] solutions, respectively. For example, the study in [18] adaptively adopts the quad-tree (QT) and binary-tree (BT) block partitions in addition to those of octrees to improve the efficiency of the coding.
II-B LiDAR Range Image Compression
Many recent works consider projecting the measured LiDAR information onto 2D RI to represent the measured distance information in a compact format. There are two types of input LiDAR information to obtain the corresponding RI s: 1) raw packet containing the LiDAR laser IDs [21], the rotation angle of the LiDAR sensors and the distance values, and 2) 3D point clouds [22, 23, 24]. Our study utilizes 3D point clouds. The obtained RI s are then intra-coded [22] or inter-coded [23, 24] in lossless and lossy manners. Here, intra-coding reduces the spatial redundancy in each RI, whereas inter-coding reduces the temporal redundancy across RI s. In R-PCC [22], which is an intra-coding method, each RI can be coded by using a lossless coding method, such as LZ4 and Deflate, to compress the floating-point format. Our study is designed for RI intra-coding and exploits the INR-based compression to represent LiDAR measurements in small storage and transmission costs.
(a) Patched depth image and mask image from LiDAR point cloud
(b) Proposed encoder and decoder
Figure 1: Overview of the proposed scheme.
II-C Implicit Neural Compression
Since the concept of INR overfits multi-dimensional signals to a small NN architecture, recent studies exploit INR architectures for image compression. Specifically, each INR architecture takes a spatial/time index of the target signals and/or the corresponding feature vector to reconstruct the corresponding attribute values, such as color information. The overfitted weights of the INR architecture are shared with the receiver’s side for signal reconstruction.
The existing INR-based compression can be classified into pixel-wise, patch-wise, and frame-wise architectures. The frame-wise architectures realized inter-coding between multiple video frames to remove temporal redundancy. They feed the frame index and/or the corresponding embeddings to the NNs to generate each frame. Neural Representations for Videos (NeRV) [25] is the first work on frame-wise video compression, and various extensions [26, 27, 28, 29, 30, 31, 32, 33, 34, 35] are proposed to improve the quality of reconstruction. However, NeRV architectures are large models when used for intra-coding each image.
The pixel-wise INR architecture [7, 36] takes the pixel index as input and reconstructs the corresponding pixel value. The patch-wise INR architecture was first proposed in [37]. Specifically, each image is divided into multiple patches, and the INR architecture takes the patch index as input to exploit the similarity of local adjacent pixels for high-quality reconstruction under the same model size. A key issue in such INR architectures is the lack of precision in high-frequency details with a small NN architecture. To represent high-frequency details under a small NN architecture, SIREN in [8] argues that sinusoidal activations work better than Rectified Linear Unit (ReLU) networks because sinusoidal activations can fit signals contained in higher-order derivatives. To address the same problem, our paper extends the training process of INRs by separating RI s into depth and mask images.
III Proposed Scheme
III-A Overview
Fig. 1 shows an end-to-end architecture of the proposed scheme. Fig. 1 (a) specifically shows the procedure to obtain RI and corresponding depth and mask images from the LiDAR 3D point cloud. We consider that the LiDAR measurement to be compressed is a 3D point cloud consisting of $N$ points, denoted as $\mathbf{P}=\{\mathbf{p}_{i}=[x_{i},y_{i},z_{i}]\ |\ i=1,\cdots,N\}$, where $x_{i},y_{i},z_{i}\in\mathbb{R}$ represent the Cartesian coordinates of the $i$-th point. The point cloud is first transformed into the spherical coordinate system. Subsequently, each point is projected onto a 2D image plane by mapping it to a pixel in a single-channel image $I\in\mathbb{R}^{W\times H}$, producing an RI. The RI is then divided into a depth image $I_{D}\in\mathbb{R}^{W\times H}$ and a mask image $I_{M}\in\{0,1\}^{W\times H}$. The depth image $I_{D}$ is further segmented into small rectangular regions, or patches, with a resolution of $\frac{W}{N_{p}}\times\frac{H}{N_{p}}$, where $N_{p}$ is the scaling factor.
Fig. 1 (b) shows the sequential operations of the encoder and decoder. In the encoder, two distinct INR s, namely the mask INR $\Phi(\cdot;\bm{\psi})$ and the depth INR $\Phi(\cdot;\bm{\omega})$ with learnable parameters $\bm{\psi}$ and $\bm{\omega}$, are trained to be overfitted to the depth image and the mask image, respectively. This training process is a pixel-wise process, which means that the parameters are trained to obtain a mapping from the pixel coordinates or patch indices on the images to their corresponding pixel values. This is achieved by sequentially providing a pair of indices to the networks. The well-trained parameters $\bm{\psi}$ and $\bm{\omega}$ are subsequently pruned and quantized as $\hat{\bm{\psi}}$ and $\hat{\bm{\omega}}$ to enhance their compactness, and we assume that these parameters are stored in storage or transmitted to content receivers as the lightweight format of the LiDAR measurements. In the decoding process, the decoder uses compressed parameters to reconstruct a mask image $\hat{I}_{M}$ and a depth image $\hat{I}_{D}$ from individual INR architectures $\Phi(\cdot;\hat{\bm{\psi}})$ and $\Phi(\cdot;\hat{\bm{\omega}})$, respectively. Similarly to the encoding process, the images are reconstructed by sequentially feeding a pair of coordinates and indices to the INR s and collecting all estimated values to form the shape of the image. We obtain the final result of RI, $\hat{I}$, by applying the mask image to the depth image to mask out any values of the pixels in the depth image corresponding to the pixels with a mask value of 1, indicating “no point”. Finally, the LiDAR point cloud is reconstructed as $\hat{\mathbf{P}}$ from $\hat{I}$ via a reverse coordinate projection process from the 2D image plane, through the spherical coordinate system, to the Cartesian coordinate system.
III-B 3D-to-2D Mapping
Our proposed method first performs a coordinate transformation for all points in the 3D point cloud $\mathbf{P}$ measured by LiDAR sensors to obtain a 2D RI $I$. Specifically, the 3D-to-2D mapping consists of two steps: 1) mapping points in the 3D Cartesian coordinate system $x$-$y$-$z$ to the spherical coordinate $\rho$-$\phi$-$\theta$, and 2) mapping points in the spherical coordinate $\rho$-$\phi$-$\theta$ to an image coordinate system $u$-$v$.
Each 3D point in the point cloud $\mathbf{p}\in\mathbf{P}$ consists of the 3D Cartesian coordinate $(x,y,z)$ first. This point is transformed into a point in the spherical coordinate $\mathbf{p}^{\prime}=(\rho,\phi,\theta)$, where $\rho,\phi,\theta$ denotes the length, pitch, and yaw of the coordinate system, as follows:
$\rho=\sqrt{x^{2}+y^{2}+z^{2}},\;\phi=\arcsin\left(\frac{z}{\rho}\right),\;\theta=\arctan\left(\frac{y}{x}\right).$ (1)
The point in the spherical coordinate is further transformed to the image coordinate $(u,v)$ to generate 2D RI $I$ as follows:
$u$ $=\left\lfloor\frac{W}{2}\times\left(\frac{\theta}{\pi}+1\right)\right\rfloor,$
$v$ $=\left\lfloor H\times\left(1-\frac{\phi+|\phi_{\mathrm{down}}|}{\phi_{\mathrm{up}}+|\phi_{\mathrm{down}}|}\right)\right\rfloor,$ (2)
[…]
Due to the sparsity of LiDAR measurements, not all pixels on the RI are guaranteed to be assigned to any 3D point. Therefore, if a pixel on $(u^{\prime},v^{\prime})$ remains unassigned after performing the 3D-to-2D mapping for all points, we set $I(u^{\prime},v^{\prime})=\rho_{\rm null}$ where $\rho_{\rm null}$ is the arbitrary value indicating that no 3D point is assigned to the pixel. In practice, $\rho_{\rm null}$ should be selected to be greater than the maximum value of $\rho$ in the LiDAR measurements or a negative value.
III-C Depth/Mask Image Construction
After 3D-to-2D mapping, the RI is then divided into a mask image $I_{M}\in\{0,1\}^{W\times H}$ and a depth image $I_{D}\in\mathbb{R}^{W\times H}$.
The mask image $I_{M}$ is to indicate whether any 3D point is assigned to each pixel on the RI or not and is defined as:
$I_{M}(u,v)=\begin{cases}1&\text{if }I(u,v)=\rho_{\rm null},\\ 0&\text{otherwise}.\end{cases}$ (3)
Given the mask image, we construct a dataset $\mathcal{D}_{M}$ for training the mask INR $\Phi(\cdot;\bm{\psi})$ which consists of pairs of the coordinates of pixels and corresponding binary values as
$\mathcal{D}_{M}=\{((u,v),I_{M}(u,v))\mid$ $u\in\{1,\ldots,W\},v\in\{1,\ldots,H\}\}.$ (4)
The depth image $I_{D}$ is the masked version of the RI. Pixels without any 3D point assignment are considered as “Do not care” ($\emptyset$) and are defined as such.
$I_{D}(u,v)=\begin{cases}\emptyset&\text{if }I(u,v)=\rho_{\rm null},\\ I(u,v)&\text{otherwise}.\end{cases}$ (5)
In addition, we divide the depth image into small rectangular areas, or patches, inspired by recent works [37] to improve the decoding performance and the quality of the reconstructed depth image. Specifically, the RI is evenly segmented into patches $I_{D}^{\prime}(i)\in\mathbb{R}^{\frac{W}{N_{p}}\times\frac{H}{N_{p}}}$, where $N_{p}$ is a scaling factor and $i=1,\cdots,N_{p}^{2}$. Each patch is assigned a patch index, allowing us to specify a pixel in the patched RI as $I_{D}^{\prime}(i,i_{u},i_{v})$, where $i$ represents the patch index and $i_{u},i_{v}$ denotes the in-patch pixel coordinates whose origin is the top left pixel in the $i$-th patch. Similarly to the mask image, we also construct a dataset $\mathcal{D}_{D}$ for training depth INR $\Psi(\cdot;\bm{\omega})$ which consists of pairs of a patch index, in-patch coordinates, and the corresponding depth values, excluding unassigned pixels, as follows:
$\mathcal{D}_{D}=\{((i,i_{u},i_{v}),I_{D}(i,i_{u},i_{v}))\mid$ $i\in\{1,\ldots,N_{p}^{2}\},$
$i_{u}\in\{1,\ldots,W/N_{p}\},$
$i_{v}\in\{1,\ldots,H/N_{p}\},$
$I_{D}(i,i_{u},i_{v})\neq\emptyset\}$ (6)
(a) Mask INR
(b) Depth INR
Figure 2: Architectures of the proposed mask and depth INRs.
III-D INR-based RI Encoder
In the encoding process, the mask INR $\Phi(\cdot;\psi)$ and the depth INR $\Phi(\cdot;\omega)$ are trained to obtain good parameters to express the coordinate-to-value relationships contained in the mask dataset $\mathcal{D}_{M}$ and $\mathcal{D}_{D}$. Figs. 2 (a) and (b) show the detailed architecture of the proposed depth INR and mask INR, respectively.
III-D1 Mask INR
Regarding the mask image, we assume the existence of a function $\Phi_{M}$, which maps each coordinate on the image to a binary value as
$\Phi_{M}:\mathbb{R}^{2}\longrightarrow\{0,1\},$ (7)
and the objective of the mask INR is to obtain parameters $\bm{\psi}$ that well approximate that mapping function as $\Phi(\cdot;\bm{\psi})\approx\Phi_{M}$ through supervised learning with the dataset $\mathcal{D}_{M}$. Fig. 2 (a) shows the detailed architecture of the proposed mask INR. The mask INR is a multi-layer perceptron (MLP) with $L$ hidden layers and $V$ nodes, and after each hidden layer, a sinusoidal function layer is used as the activation function. The network sequentially receives the coordinate $(u,v)$ from the mask dataset $\mathcal{D}_{M}$ and regresses a binary value as its output. Regression loss is computed using binary cross entropy (BCE) loss function between all output values $\Phi((u,v);\bm{\psi})$ and the corresponding true values $I_{M}(u,v)$ as follows:
$\mathcal{L}_{\mathrm{BCE}}(\bm{\psi})=-\frac{1}{HW}\sum_{u}^{H}\sum_{v}^{W}[I_{M}(u,v)\log\left(\Phi((u,v);\bm{\psi})\right)$
$+(1-I_{M}(u,v))\log\left(1-\Phi((u,v);\bm{\psi})\right)].$ (8)
III-D2 Depth INR
Similar to the mask INR, we also assume the existence of another function $\Phi_{D}$ regarding the depth image, which maps the pair of a patch index and an in-patch pixel coordinate to a depth value as
$\Phi_{D}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{1}.$ (9)
and the objective of the depth INR is to obtain parameters $\bm{\omega}$ for good approximation of $\Phi_{D}$ as $\Phi((i,i_{u},i_{v});\bm{\omega})\approx\Phi_{D}$. Figs. 2 (b) shows the detailed architecture of the proposed depth INR. The depth INR shares its structure with the mask INR, with the exception of the input layer, which accepts 3 values. In the training process, pairs of patch index $i$ and an in-patch coordinate $(i_{u},i_{v})$ are sequentially passed to the depth INR network from the depth dataset $\mathcal{D}_{D}$, and the corresponding depth values are regressed. We employ mean squared error (MSE) loss as a regression loss function for the depth INR as
$\mathcal{L}_{\mathrm{MSE}}(\bm{\omega})=\frac{1}{HW}\sum_{i}^{N_{p}^{2}}\sum_{i_{u}}^{W/N_{p}}\sum_{i_{v}}^{H/N_{p}}\|\Phi((i,i_{u},i_{v});\bm{\omega})-I_{D}(i,i_{u},i_{v})\|^{2}.$ (10)
III-E Model Compression
Parameters $\bm{\psi}$ and $\bm{\omega}$ become effectively compressed representations of the depth and mask images after a thorough training. We introduce a series of parameter compression processes for both to further improve their compactness.
III-E1 Model Pruning
As an initial step in our parameter compression procedure, we implement global unstructured pruning for parameters in both depth and mask INR s. Given a threshold $w_{q}$ for the magnitude of parameters, each parameter $w$ is determined to be retained or pruned based on the following criteria:
$\hat{\bm{\omega}}=\begin{cases}\bm{\omega}&\bm{\omega}\geq\bm{\omega}_{q},\\ \mathbf{0}&\mathrm{otherwise}.\end{cases}$ (11)
To guarantee that the pruned parameters are of good expression, we subsequently retrain the parameters to fine-tune using the same dataset $\mathcal{D}_{M}$ and $\mathcal{D}_{D}$.
III-E2 Model Quantization and Encoding
The pruned and fine-tuned parameters are uniformly quantized to a bit depth of $N_{b}$. This quantization is layer-wise, meaning that given a parameter set corresponding to each layer in the depth and mask INRs as $\bm{\mu}\in\hat{\bm{\omega}}$, a quantized parameter set $\bm{\mu}_{q}$ is obtained as follows:
$\bm{\mu}_{q}=\mathrm{round}\left(\frac{\bm{\mu}-\bm{\mu}_{\min}}{2^{N_{b}}}\right)s+\bm{\mu}_{\min},\;s=\frac{\bm{\mu}_{\max}-\bm{\mu}_{\min}}{2^{N_{b}}},$ (12)
where $\mathrm{round}(\cdot)$ is a rounding function to the nearest integer and $\bm{\mu}_{\max}$ and $\bm{\mu}_{\min}$ are the maximum and minimum values in $\bm{\mu}$. The quantized tensor $\bm{\mu}_{q}$ is finally coded into a binary sequence using Huffman coding. It is noteworthy that the quantized parameters $\bm{\mu}_{q}$ are likely to assume values near zero, particularly for smaller bit depths. Consequently, Huffman coding demonstrates its effectiveness in reducing the overall size of encoded parameters.
III-F RI Decoder
The decoding process of RI is a simple feedforward process involving the mask INR and depth INR with optimized parameters $\hat{\bm{\psi}}$ and $\hat{\bm{\omega}}$. The mask image is reconstructed by feeding the coordinate sets $\{(u,v)\ |\ u\in\{1,\cdots,W\},v\in\{1,\cdots,H\}\}$ to the mask INR. The resulting binary values are then reshaped to construct the $W\times H$ mask image $\hat{I}_{M}$.
The depth image reconstruction is in a two-stage manner. Each patch is first reconstructed by feeding the sets of pairs of a patch index and an in-patch coordinate $\{(i,i_{u},i_{v})\ |\ i\in\{1,\cdots,N_{p}^{2}\},i_{u}\in\{1,\cdots,W\},i_{v}\in\{1,\cdots,H\}\}$ to the depth INR The reconstructed patches are then gathered to build the complete depth image $\hat{I}_{D}$. Finally, the reconstructed RI $\hat{I}$ is obtained as
$\hat{I}(u,v)=\begin{cases}\hat{I}_{D}(u,v)&\hat{I}_{M}(u,v)=0,\\ \rho_{\rm null}&\mathrm{otherwise}.\end{cases}$ (13)
III-G 2D-to-3D Mapping
The concluding phase of our decoding procedure is the reconstruction of a 3D point cloud through a 2D-to-3D mapping against the reconstructed RI. When the pixel $(u,v)$ on the RI has a valid point depth, i.e., is not $\rho_{\rm null}$, the corresponding point in the spherical coordinate $\hat{\mathbf{p}}^{\prime}=(\hat{\rho},\hat{\phi},\hat{\theta})$ is obtained as follows:
$\hat{\rho}$ $=\hat{I}(u,v),$
$\hat{\phi}$ $=\left(1-\frac{v}{H}\right)\left(\phi_{up}+|\phi_{\text{down}}|\right)-|\phi_{\text{down}}|,$
$\hat{\theta}$ $=-\left(2\frac{u}{W}-1\right)\pi.$ (14)
Finally, the 3D points in the spherical coordinate are transformed into the 3D Cartesian coordinate $\hat{\mathbf{p}}=(\hat{x},\hat{y},\hat{z})$ as
$\hat{x}=\hat{\rho}\cos\hat{\phi}\cos\hat{\theta},\;\hat{y}=\hat{\rho}\cos\hat{\phi}\sin\hat{\theta}\;,\hat{z}=\hat{\rho}\sin\hat{\phi}.$ (15)
IV Experiments
IV-A Settings
Dataset: We use the KITTI dataset [10] as our source of 3D point cloud data. For R-D performance, we evaluate the KITTI Odometry dataset, using frames 00, 25, 50, 75, and 100 from sequences 00 to 06. For downstream tasks, we use the KITTI 3D Object Detection dataset, specifically evaluating frame 000134 to analyze the impact of compression on object detection accuracy. All data were collected with a Velodyne 64 scanner that features 64 laser scan lines and an azimuth resolution of 0.09 degrees.
Figure 3: Chamfer distance as a function of bitrates across different sequences of KITTI’s LiDAR point cloud. From left to right: (left) Seq:00-00, (middle) Seq:00-25, and (right) Seq:00-50.
Refer to caption (a) Seq:00-00 Refer to caption (b) JPEG
bpp:2.89
CD:0.717
Refer to caption (c) JPEG2000
bpp:2.55
CD:0.436
Refer to caption (d) HEIF
bpp:2.35
CD:0.339
Refer to caption (e) AVIF
bpp:1.41
CD:0.324
Refer to caption (f) COIN
bpp:3.02
CD:0.991
Refer to caption (g) G-PCC
bpp:2.62
CD:0.137
Refer to caption (h) Draco
bpp:2.81
CD:0.127
Refer to caption (i) R-PCC (Deflate)
bpp:2.73
CD:0.0867
Refer to caption (j) R-PCC (LZ4)
bpp:3.63
CD:0.229
Refer to caption (k) Proposed
(pixel-wise)
bpp:2.99
CD:0.123
Refer to caption (l) Proposed
bpp:2.55
CD:0.0902
Refer to caption (m) Seq:00-25 Refer to caption (n) JPEG
bpp:2.47
CD:1.01
Refer to caption (o) JPEG2000
bpp:2.55
CD:0.441
Refer to caption (p) HEIF
bpp:2.67
CD:0.323
Refer to caption (q) AVIF
bpp:1.51
CD:0.308
Refer to caption (r) COIN
bpp:3.02
CD:0.957
Refer to caption (s) G-PCC
bpp:2.30
CD:0.136
Refer to caption (t) Draco
bpp:2.53
CD:0.128
Refer to caption (u) R-PCC (Deflate)
bpp:2.48
CD:0.106
Refer to caption (v) R-PCC (LZ4)
bpp:4.05
CD:0.235
Refer to caption (w) Proposed
(pixel-wise)
bpp:2.04
CD:0.107
Refer to caption (x) Proposed
bpp:2.55
CD:0.0982
Figure 4: Snapshots of the reconstructed LiDAR point clouds in proposed and baseline methods. Here, (b)-(l) and (n)-(x) show the reconstructed point clouds of sequences 00-00 and 00-25, respectively.
TABLE I: The list of BD-CD $\uparrow$ for KITTI dataset for the different sequences.
Seq. JPEG JPEG2000 HEIF AVIF COIN G-PCC Draco
R-PCC
R-PCC
00-00 0.887 0.426 0.266 0.149 0.955 0.054 0.054 0.012 0.124
00-25 0.924 0.435 0.263 0.137 0.894 0.029 0.040 0.007 0.146
00-50 1.198 0.531 0.294 0.166 1.050 0.024 0.025 0.014 0.156
00-75 1.261 0.563 0.309 0.169 0.905 0.022 0.044 -0.019 0.091
00-100 0.806 0.383 0.232 0.130 1.019 0.036 0.040 -0.026 0.096
01-00 0.825 0.391 0.242 0.109 0.990 0.041 0.064 -0.039 0.101
01-25 0.753 0.385 0.250 0.142 0.999 0.100 0.113 -0.017 0.114
01-50 0.822 0.393 0.259 0.147 1.013 0.065 0.086 0.000 0.120
01-75 1.072 0.476 0.287 0.153 0.897 0.053 0.060 -0.004 0.102
01-100 1.037 0.485 0.281 0.170 1.001 0.045 0.074 0.014 0.156
02-00 0.996 0.441 0.259 0.121 0.902 0.015 0.032 -0.044 0.137
02-25 0.769 0.380 0.236 0.129 0.919 0.036 0.074 -0.011 0.110
02-50 0.775 0.384 0.234 0.143 0.933 0.067 0.082 -0.024 0.094
02-75 0.743 0.369 0.249 0.149 0.882 0.036 0.095 -0.022 0.085
02-100 0.838 0.389 0.250 0.170 1.063 0.057 0.096 -0.008 0.095
03-00 1.065 0.440 0.251 0.139 0.927 0.021 0.021 0.012 0.162
03-25 1.241 0.592 0.319 0.118 0.873 -0.006 -0.012 -0.026 0.106
03-50 0.905 0.409 0.244 0.125 0.975 0.013 0.017 -0.052 0.092
03-75 0.816 0.364 0.219 0.109 0.856 0.028 0.034 -0.067 0.054
03-100 0.810 0.369 0.227 0.122 1.025 0.040 0.067 -0.054 0.077
04-00 0.769 0.352 0.215 0.127 0.987 0.031 0.036 -0.027 0.159
04-25 1.037 0.502 0.265 0.083 0.954 0.006 0.013 -0.042 0.105
04-50 1.103 0.476 0.272 0.147 1.008 0.024 0.047 -0.006 0.121
04-75 0.919 0.401 0.232 0.131 0.858 0.005 0.011 -0.038 0.104
04-100 0.954 0.435 0.263 0.162 1.079 0.047 0.075 -0.003 0.102
05-00 0.807 0.402 0.240 0.144 0.942 0.036 0.038 0.008 0.149
05-25 0.811 0.396 0.243 0.168 0.929 0.042 0.052 -0.024 0.081
05-50 1.271 0.576 0.300 0.131 0.920 0.001 0.018 -0.037 0.089
05-75 1.380 0.634 0.333 0.118 0.852 -0.013 -0.014 -0.043 0.106
05-100 1.064 0.467 0.279 0.170 1.007 0.017 0.039 -0.030 0.084
06-00 0.836 0.388 0.219 0.093 0.868 0.010 0.003 -0.034 0.146
06-25 1.085 0.477 0.246 0.073 0.864 -0.001 0.001 -0.058 0.094
06-50 1.015 0.459 0.238 0.076 0.871 -0.007 -0.020 -0.064 0.080
06-75 1.256 0.612 0.321 0.078 0.830 -0.013 -0.013 -0.046 0.099
06-100 0.793 0.374 0.213 0.075 0.963 0.025 0.034 -0.065 0.080
Average 0.961 0.444 0.259 0.131 0.943 0.028 0.041 -0.025 0.109
TABLE II: 3D object detection accuracy $\uparrow$ for different model sizes.
IoU G-PCC R-PCC Proposed
Type Object Low Mid High Low Mid High Low Mid High
2D IoU Pedestrian 0.07 0.27 0.68 0.08 0.16 0.18 0.43 0.45 0.52
Cyclist 0.22 0.70 0.83 0.30 0.55 0.58 0.86 0.90 0.82
Car - 0.19 0.20 0.30 0.26 0.22 0.22 0.23 0.26
BEV IoU Pedestrian - 0.04 0.31 - 0.04 0.02 0.74 0.67 0.57
Cyclist - 0.15 0.81 0.03 0.15 0.16 0.72 0.71 0.73
Car - - - - - - - - -
3D IoU Pedestrian - 0.04 0.28 - 0.04 0.02 0.54 0.51 0.50
Cyclist - 0.15 0.75 0.02 0.142 0.142 0.64 0.68 0.67
Car - - - - - - - - -
TABLE III: Average decoding latency $\downarrow$
Method Latency per sequence (ms)
JPEG 0.49
JPEG2000 0.54
HEIF 0.51
AVIF 0.48
COIN 0.71
G-PCC 2.00
Draco 3.00
R-PCC (Deflate) 11.5
R-PCC (LZ4) 10.3
Proposed 0.69
Metric: Regarding the metrics for the decoded 3D point clouds, we follow the common practice in the community using chamfer distance (CD). CD is defined as:
$\mathrm{CD}=\frac{1}{2}\bigg{\{}\frac{1}{|\mathbf{P}|}\sum_{\bm{p}\in\mathbf{P}}\min_{\widehat{\bm{p}}\in\widehat{\mathbf{P}}}\|\bm{p}-\widehat{\bm{p}}\|_{2}+\frac{1}{|\widehat{\mathbf{P}}|}\sum_{\bm{\widehat{p}}\in\widehat{\mathbf{P}}}\min_{\bm{p}\in\mathbf{P}}\|\bm{p}-\widehat{\bm{p}}\|_{2}\bigg{\}},$ (16)
where $\mathbf{P}$ is the set of 3D points in the original point cloud and $\widehat{\mathbf{P}}$ is the set of 3D points in the decoded point set.
For the R-D performance assessment between the proposed method and the baselines, we use the Bjøntegaard delta chamfer distance (BD-CD) [38] for calculating average CD improvement between R-D curves for the same bitrate, where positive values denote CD improvement compared to the baselines.
Network Architecture Details: Both INR architectures are designed to effectively approximate the coordinate-to-value mappings in the mask and depth images derived from RI. The mask INR is an MLP with a fixed depth of $L=6$ layers. We experimented with varying the number of nodes $V$ in each hidden layer to evaluate the impact on performance and compression efficiency. The values of $V$ considered are $\{10,19,24,28,31,34,37,40\}$.
The network takes as input the pixel coordinates $(u,v)\in\mathbb{R}^{2}$ from the mask dataset $\mathcal{D}_{M}$ and outputs a scalar value representing the mask at that coordinate. The architecture is structured as follows:
  • Input Layer: The coordinates of the pixels $(u,v)$.
  • Hidden Layers: It consists of $L=6$ hidden layers, each with $V$ nodes. Each hidden layer employs the sinusoidal activation function to introduce periodicity and enable the network to model high-frequency variations in the mask image.
  • Output Layer: A single node with the sigmoid activation function produces an output in the range $(0,1)$, suitable for binary classification of mask values.
The Depth INR is also implemented as an MLP with a fixed depth of $L=6$ layers. We also consider the different number of nodes $V$ in each hidden layer, choosing $\{28,31,34,37,40,42,45\}$ to evaluate the trade-off between model capacity and compression.
The input to the depth INR is a concatenation of the patch index $i$ and the in-patch pixel coordinates $(i_{u},i_{v})\in\mathbb{R}^{2}$, resulting in a 3-dimensional input vector. Since we set the patch scaling factor to $N_{p}=16$, the patch index $i$ ranges from 00 to $255$. The architecture of the depth INR is as follows:
  • Input Layer: The concatenated input vector $(i,i_{u},i_{v})$.
  • Hidden Layers: It contains $L=6$ hidden layers, each with $V$ nodes. Similarly to the mask INR, the sinusoidal activation function is applied to each hidden layer to capture the complex variations in the depth image.
  • Output Layer: A single node with a linear activation function (identity function) to output the estimated depth value $\hat{I}_{D}(i,i_{u},i_{v})\in\mathbb{R}$.
Hyperparameter Details: We use separate hyperparameter settings for mask and depth INRs. The general settings for both INRs include the Adam optimizer, an initial learning rate of $1\times 10^{-3}$, 3,000 training epochs, and a batch size of 1. For depth INR, we adopt the cosine annealing scheduler with a warmup phase. The initial learning rate is set to $1\times 10^{-4}$, and the warmup period lasts for 300 epochs. The minimum learning rate is set to $1\times 10^{-12}$.
Model Compression Details: A global unstructured pruning is used for model pruning. The pruning ratio (sparsity) was varied from 0 to 1 to adjust the sparsity of the model parameters. For each pruning ratio, we determined the corresponding threshold $\omega_{q}$ to control which parameters were pruned. A higher pruning ratio results in more parameters being set to zero. After pruning, we fine-tuned the model using the same dataset to recover any potential performance loss.
To further compress the pruned and fine-tuned model, We perform uniform quantization to the parameters. The quantization bit depth $N_{b}$ was varied from 4 to 32 bits to balance compression performance and model precision. The quantized parameters were then encoded using Huffman coding to further reduce the model size.
Baselines: We evaluate our proposed method by comparing it with existing baselines in both geometric 3D point cloud compression and 2D image compression.
  1. 1)
    As baselines for 3D point cloud compression, we select G-PCC within the PCC family. We refer to the MPEG reference implementation TMC13-v14.0 for octree geometry compression.
  2. 2)
    We also select Draco [39] as the 3D point cloud compression baseline which also belongs to the PCC family. We use the official implementation of the Draco encoder that performs KD-tree-based compression [40].
  3. 3)
    As conventional image compression baselines, we select JPEG, JPEG2000, High-Efficiency Image File Format (HEIF), and AV1 Image File Format (AVIF). We convert floating-point valued RI s into 8-bit precision in advance when using these methods.
  4. 4)
    R-PCC [22] is an RI based LiDAR compression baseline. It maps LiDAR point clouds to RI s and performs intra-coding using floating-point lossless coding methods. Here, we use LZ4 and Deflate for coding methods due to their fast decompression.
  5. 5)
    COmpression with Implicit Neural representations (COIN) [7] is an INR–based image compression baseline. The INR architecture is trained to obtain a direct mapping of the pixel coordinate to each pixel value of RI. We assume that COIN serves as a reliable indicator of the efficiency of our depth/mask separation strategy, as it does not employ the process.
Implementation Detail: All the evaluations exhibited in this paper are performed with CPUs of Intel Core i9-10850K and i9-13900KF and with GPUs of NVIDIA GeForce RTX 3080 and 4070. NNs for COIN and our proposed method are implemented, trained, and evaluated using PyTorch 2.2.0 with Python 3.10.
IV-B Comparison with Baselines
IV-B1 Rate-Distortion Performance
We show the R-D performance of our proposed method and baselines. Figs. 3 show the CD between the original and reconstructed LiDAR point clouds against various bitrates, i.e., bit per point (bpp). We observe the following findings:
  • The proposed method achieves the best 3D reconstruction quality in the bpp range of 1 through 2.5.
  • Image compression methods suffer from quality saturation due to the precision disparity between RI and the typical 8-bit precision image.
  • PCC methods do not have saturation since they compress the geometry information with 10-bit precision.
  • In R-PCC, R-D performance highly depends on the lossless coding method.
  • The INR-based method requires a large model size for reconstructing high-quality RI.
Figs. 4 (a)-(x) show the snapshots of the original and reconstructed LiDAR point clouds in each method at approximately 2.55 bpp regimes. Here, Figs. 4 (a)-(l) and (m)-(x) use the sequences 00-00 and 00-25, respectively. The proposed method can reconstruct a clean point cloud at the same bitrate. However, all PCC, image compression, and INR-based compression methods contain circular noises and/or decrease the number of 3D points in the reconstructed LiDAR point clouds. A circular noise still remains in R-PCC methods as well.
Table I lists the BD-CD of the proposed method against the baselines in each sequence of LiDAR point clouds. It shows that the proposed method achieves the best 3D reconstruction quality in the same bitrate range against PCC, image compression, and INR-based compression methods irrespective of LiDAR sequences. In addition, the performance gap between R-PCC and proposed methods depends on the lossless coding method. When R-PCC uses a low-efficiency coding method, such as LZ4, for fast decoding, the proposed scheme achieves better R-D performance than R-PCC.
IV-B2 Downstream Task
We then discuss the impact of our RI compression method on the performance of downstream tasks on the LiDAR point cloud. We selected 3D object detection as a representative example of downstream tasks. We used PointPillar [41] as a 3D object detector, and we used our method and G-PCC and R-PCC (Deflate), which are baselines with better BD-CD performance, to compress the point clouds and fed the reconstructed ones into PointPillar. In addition, we assumed three model sizes (approximately bpp 1.5 as Low, 2.5 as Middle, and 3.5 as High) in each compression method. Table III shows the 2D Intersection over Union (IoU), Birds Eye View (BEV) IoU, and 3D IoU in each object for each compression approach. Especially in limited bpp scenarios, the proposed approach performs better than other baselines in most IoU metrics.
IV-B3 Decoding Latency
Table III shows the average decoding latency of the proposed and baseline methods for LiDAR sequence of 00-00. The decoding latency values for the proposed method and the baselines are the total time required from RI decoding to the 2D-to-3D mapping. The decoding delay of the proposed method is comparable to that of image compression methods and has more than 65.5% and 93.3% reduction compared to the PCC and R-PCC methods, respectively. This means that the proposed method approaches the decoding delay of image compression methods and achieves 3D reconstruction quality comparable to PCC/R-PCC methods.
(a) R-D performance.
(b) Convergence speed.
Figure 5: Patch-wise vs. pixel-wise.
(a) CD for different pruning sparsity.
(b) Quantization with different $N_{b}$.
Figure 6: Model compression performance.
IV-C Ablation Study
IV-C1 Impact of Patch-wise INR Architecture
The proposed depth INR exploits the patch-wise architecture, whereas the pixel-wise architecture can be used for the depth INR. Fig. 6 (a) shows the CD of the proposed patch-wise INR and pixel-wise INR architectures as a function of bitrates under the different sequences of KITTI’s LiDAR point cloud. We can see that the patch-wise depth INR achieves better CD than the pixel-wise architecture at large bitrate regimes in every LiDAR sequence. Specifically, BD-CD between the patch-wise and pixel-wise architectures is 0.032, 0.017, and 0.011 in the sequences 00-00, 00-25, and 00-50, respectively. The effects on the visual quality are shown in Figs. 4 (k), (l), (w), and (x), respectively.
Fig. 6 (b) shows the CD performance of INR-based image compression methods as a function of the learning epochs. Our patch-wise architecture boosts the convergence speed by up to 13$\times$ compared to the pixel-wise architecture, and the fast convergence results in a short encoding delay.
IV-C2 Impact of Model Compression
After the encoder trains the depth and mask INR architectures, the trained weights are pruned and quantized for compression. Here, the proposed method can set different pruning ratios and bit depths for the depth and mask INRs. This section discusses the impact of model pruning and quantization on both INR model compression.
Figs. 6 (a) and (b) show the effect of model pruning and quantization for the depth and mask architectures. In pruning, the mask INR is similar in performance to the full model, although the sparsity is approximately 70%. However, pruning the model for the depth INR causes quality degradation even though the sparsity is only 10%. For quantization, a 16-bit model still retains almost the same CD as the original 32-bit model in depth INR, while the mask INR can reduce the number of bits to 11.
IV-C3 Impact of Patch Size
Figure 7: Chamfer distance under the different patch sizes. Here, LiDAR sequence is 00-00.
Refer to caption (a) Seq:00-00 Refer to caption (b) $N_{p}$: 2
Refer to caption (c) $N_{p}$: 4 Refer to caption (d) $N_{p}$: 8
Refer to caption (e) $N_{p}$: 16 Refer to caption (f) $N_{p}$: 32
Refer to caption (g) $N_{p}$: 64
Figure 8: Snapshots of the reconstructed LiDAR point clouds in proposed methods under the different patch sizes $N_{p}\times N_{p}$. Here, (b)-(g) show the reconstructed point clouds of sequence 00-00.
The proposed depth INR uses the patch-wise architecture, and thus the depth image is divided into patches of size $N_{p}\times N_{p}$. We set the patch size to $16\times 16$ for best performance. This section discusses an effect of the patch size on the quality of the reconstructed point cloud.
Fig. 7 shows CD performance varying $N_{p}$, and Fig. 8 demonstrates the corresponding snapshots of the reconstructed point cloud. We consider all the variants using the same model size in Fig. 6 with a sparsity of 0.0 and $N_{b}$ of 32. As mentioned, $N_{p}=16$ is the best performance. For example, in smaller patch sizes, such as ($N_{p}=2$ and $N_{p}=4$), the patches cover larger areas, resulting in a loss of fine local detail and reducing the model’s ability to capture intricate depth variations.
V Conclusion and Future Work
We proposed a novel RI-based LiDAR point cloud compression method. The proposed method is designed to efficiently compress floating-point RIs using INR-based techniques and features a sophisticated architecture that combines separated learning for mask and depth images, patch-wise learning for depth images, and model compression. Experiments on the KITTI dataset show that the proposed method improves 3D reconstruction quality at low bit rates compared to existing image, point cloud, RI, and INR-based compression methods, resulting in a BD-CD improvement of up to 0.961, and the improvement also brings better performance in the downstream task of 3D object detection.
The proposed method has two limitations: encoding delay and transformation loss from RIs to 3D point clouds. While existing baselines require only a few milliseconds for encoding, implicit neural compression, including the proposed method, takes from tens of minutes to several hours. In summary, the proposed method’s long encoding delay and short decoding delay make it well-suited for offline applications of LiDAR point clouds. In addition, RIs with limited spatial resolution will lead to irreversible point loss during 2D-to-3D decoding, potentially degrading the performance of downstream tasks. In future work, we will consider integrating a point cloud generator [42] to obtain a denser point cloud from the limited resolution of RIs.